OpenStudy (anonymous):

Find the points of the graph of y= x^2 + 4ln(x) at which the tangent line is parallel to the line y= 6x - 3. Thanks!

OpenStudy (anonymous):

the solpe of the tangent must be equal to the slope of the line which is 6

OpenStudy (anonymous):

here dy/dx =2x + 4/x by the problem 2x +4/x =6 or x+2/x=3 2=(3-x)x or x^2 -3x+2=0 or (x-1)(x-2)=0 thus x=1,2 thus at x=1 or 2 tangent to the given curve is parallel to the given line when x=1 y = 1+4ln1=1 when x=2 y=4+4ln2 thus the reqd pts are (1,1) and (2,4+4ln2)